According to a student conducted by the grade 12 students P155 is the average monthly expense for cellphone loads of high school students in their province. A statistics student claims that this amount has increased since January of this year. do you think his claim is acceptable if a random sample of 50 students has an average monthly expense of P165 for cellphone loads? Using 5% level of significance, assume that a population standard deviation is P52

1
Expert's answer
2022-06-01T10:27:54-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le155"

"H_1:\mu>150"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c =1.6449."

The rejection region for this right-tailed test is "R = \{z:z> 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{165-155}{52/\sqrt{50}}=1.3598"


Since it is observed that "z=1.3598<1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed is "p=P(Z>1.3598)=0.086947," and since "p=0.086947>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 155, at the "\alpha = 0.05" significance level.